# Exploring the Adversarial Frontier: Quantifying Robustness via Adversarial Hypervolume

### v1

@@ III-B Adversarial Frontier
Drawing on the multiobjective adversarial framework outlined in Section III-A, we incorporate the multiobjective optimization principles to pinpoint optimal trade-off solutions. We begin by elucidating the concepts of Pareto dominance and Pareto optimality, which are central to multi-objective optimization and represent the most efficient trade-off solutions. Subsequently, we employ an example to demonstrate the adversarial frontier delineated by these principles and to showcase the adversarial frontiers of various defensive models.
@@
Figure 2: The adversarial frontiers of different models computed using a randomly selected image from CIFAR-10 test set. R is for ResNet and WR is for WideResNet. The solid line means the example is correctly classified, the dashed line means the example is misclassified.
To give an intuitive understanding of the adversarial frontier, we first implement the confidence loss and distance loss and then visualize the adversarial frontiers of several defensive models. Specifically, we employ the $\ell_{2}$ norm as the metric for distance loss, owing to its widespread acceptance and inherent simplicity. Subsequently, we implement the confidence loss in the following form:
$\mathcal{L}_{\text{CE\_conf}}(f,x,y,\delta)$ =$=\left\{\begin{array}[]{cc}\text{CE}_{\text{norm}}&\text{if }\text{CE}_{\text{norm}}>0\\ 0&\text{otherwise}\end{array}\right.,$ (2)
$\text{CE}_{\text{norm}}(f,x,y,\delta)$ $=\frac{\log m+\log f_{y}(x+\delta)}{\log m}.$
This expression denotes the normalized cross-entropy loss, selected for its characteristic whereby a negative value definitively signifies an instance misclassified by the model. A detailed discussion of this attribute is provided in Section VI-A.
Upon resolving the optimization problem in Equation (1), we obtain the adversarial frontiers corresponding to various models [^1] Here, we simply optimize the confidence loss using established attacks (PGD) with multiple fixed perturbations to derive these solutions, which is considered a basic approach to addressing the multiobjective problem (point-by-point).. Specifically, we examine three adversarially trained models on the CIFAR-10 dataset [42]: ResNet-18 (R-18) [43], WideResNet-70-16 (WR-70-16) [41], and WideResNet-34-10 (WR-34-10) [44], all sourced from the RobustBench repository [16]. Each model has been trained with an $\ell_{2}$ perturbation budget of $\epsilon=0.5$. For contrastive analysis, we incorporate a standardly trained WideResNet-28-10 model. We plot the adversarial frontiers of these models using a randomly selected image from the test set in Figure 2.
Figure 2 illustrates that the normalized confidence loss of the standardly trained model decreases significantly with minimal perturbation, in contrast with the adversarially trained counterparts that preserve the confidence level within a small perturbation margin. Nonetheless, it is challenging to compare the robustness of models such as R-18 and WR-70-16 with reference to the adversarial frontiers derived from a singular image. To address this, we introduce the adversarial hypervolume indicator, a new metric designed to quantify the robustness of models uniformly across different perturbation levels.
@@ III-C Adversarial Hypervolume
III-D Adversarial Hypervolume and Adversarial Accuracy
In this section, we demonstrate that our proposed adversarial hypervolume metric delivers more granular insights than adversarial accuracy. Drawing on a multiobjective framework, we propose the adversarial frontier curve concept, and we clarify the association between the adversarial hypervolume and the area under the curve. Furthermore, our analysis examines the interconnections among the adversarial hypervolume metric, adversarial accuracy, and adversarial sparsity.
Marginal Confidence Loss. Within our multiobjective adversarial framework, we can implement a variety of confidence loss functions. We have chosen to utilize the marginal confidence loss to examine the correlation between our proposed metric and adversarial accuracy in this section. The marginal confidence loss is mathematically expressed as:
Compared to normalized cross-entropy loss $\text{CE}_{\text{norm}}$, the marginal confidence loss demonstrates increased sensitivity to errors in classification. A negative MAR indicates a model’s misclassification, satisfying both the necessary and sufficient conditions for a classification error.
@@ III-D Adversarial Hypervolume and Adversarial Accuracy
Adversarial Hypervolume. The adversarial hypervolume, denoted by $\text{AH}_{\text{MAR}}(f,x,y,\epsilon)$, where $\epsilon$ represents the perturbation budget, is calculated by integrating the adversarial frontier curve across the full spectrum of possible perturbations. The relationship is formally expressed as follows:
@@
where $\mathbf{e}$ denotes the error term that quantifies the discrepancy between the integral and the computed adversarial hypervolume. Furthermore, we provide a convergence analysis of the error term $\mathbf{e}$ in Section III-E.

### v2

@@ III-B Adversarial Frontier
Drawing on the multiobjective adversarial framework outlined in Section III-A, we incorporate the multiobjective optimization principles to determine optimal trade-off solutions. We begin by presenting the concepts of Pareto dominance and Pareto optimality, which are central to multi-objective optimization and represent the most efficient trade-off solutions. Subsequently, we employ an example to demonstrate the adversarial frontier delineated by these principles and to showcase the adversarial frontiers of various defensive models.
@@
Figure 2: An example of adversarial frontiers. The adversarial frontiers of different models computed using a randomly selected image from CIFAR-10 test set. R is for ResNet and WR is for WideResNet. The label represents the name of the model in the RobustBench library.
Upon resolving the optimization problem in Equation (2) implemented with the marginal confidence loss in Equation (3) and $\ell_{2}$ norm, we obtain the adversarial frontiers corresponding to various models [^1] Here, we simply optimize the confidence loss using established attacks (PGD) with multiple fixed perturbations to derive these solutions, which is considered a basic approach to addressing the multiobjective problem (point-by-point).. Specifically, we examine three adversarially trained models on the CIFAR-10 dataset [53]: ResNet-18 (R-18) [54, 55], and WideResNet-34-10 (WR-34-10) [56], all sourced from the RobustBench repository [16]. Each model has been trained with an $\ell_{2}$ perturbation budget of $\epsilon=0.5$. For contrastive analysis, we incorporate a standardly trained WideResNet-28-10 model. We plot the adversarial frontiers using a randomly selected image from the test set in Figure 2.
Figure 2 illustrates that the normalized confidence loss of the standardly trained model decreases significantly with minimal perturbation, in contrast with the adversarially trained counterparts that preserve the confidence level within a small perturbation margin. Nonetheless, it is inadequate to rely solely on the adversarial accuracy of a certain perturbation level to compare the robustness of models (e.g., R-18 [54, 55]). To address this, we introduce the adversarial hypervolume indicator, a new metric designed to quantify the robustness of models uniformly across different perturbation levels.
@@ III-C Adversarial Hypervolume
III-D Convergence Analysis
In this section, we explored to measure an averaged version of robustness by integrating the adversarial frontier curve across the full spectrum of possible perturbations. While this integral is a direct representation of robustness, it is challenging to compute in practice as it requires the adversarial frontier curve to be known. Thus, we approximate the integral using a set of discrete points on the adversarial frontier curve, presenting the adversarial hypervolume as an approximation of the integral. We provide a convergence analysis of the error term between the adversarial hypervolume and the integral of the adversarial frontier in this section.
@@ III-D Convergence Analysis
Integral of Adversarial Frontier. Averaged version of robustness can be obtained by integrating the adversarial frontier curve across the full spectrum of possible perturbations. This integral is represented as:
@@
where $\mathbf{e}$ denotes the error term that quantifies the discrepancy between the integral and the computed adversarial hypervolume.
Approximation Error. For an image $(x,y)$, a classifier $f$, its corresponding adversarial frontier as $\mathcal{L}_{\text{conf}}=F(z)$, where $z$ is the magnitude of the perturbation, usually measured by the $p$-norm. It must be a monotonic non-increasing function by definition of the adversarial front. Moreover, we assume that the following conditions hold:
Assumption 1.
[…]
$=\frac{\epsilon^{2}}{N^{2}}L_{0}$
$\mathbf{e}=N\mathbf{e}_{i}\leq\frac{\epsilon^{2}}{N}L_{0}$ (8)
IV Algorithm
This section delineates algorithms designed to identify adversarial frontiers for adversarial training approach that seeks to augment model robustness over a range of perturbation levels. Furthermore, this section presents the implementation details and analyzes the complexity inherent in these algorithms.
IV-A Adversarial Hypervolume Computation
Algorithm 1 Adversarial Hypervolume Computation
1:  Input: Data point $(x,y)$, perturbation budget $\epsilon$, model $f$, evaluation points $N$
2:  Output: Adversarial Hypervolume AH
3:  $\text{AF}=\emptyset$
4:  for $i=1$ to $N$ do
5:     $\delta=\text{PGD}(f,x,y,\frac{\epsilon}{N})$
6:     if $\mathcal{L}_{\text{conf}}(f,x,y,\delta)<0$ then
7:        $\text{AF}=\text{AF}\cup\{(\frac{i}{N},0)\}$
8:        break
9:     else
10:        $\text{AF}=\text{AF}\cup\{(\frac{i}{N},\mathcal{L}_{\text{conf}}(f,x,y,\delta))\}$
11:     end if
12:  end for
[…]
Calculating the Hypervolume. To accurately calculate the hypervolume enclosed by the adversarial frontier, established algorithms from multiobjective optimization literature can be employed [57, 58]. Specifically, in the case of a two-objective problem, the adversarial frontier is first organized in ascending order according to the perturbation level values. Subsequently, the volume within each interval $[\frac{i\epsilon}{N},\frac{(i+1)\epsilon}{N}]$ is determined by calculating the product of the interval’s length and the maximum objective value at the end of the interval.
Batched Adversarial Hypervolume. Implementing batch processing significantly enhances the computational efficiency of neural networks. In particular, the PGD algorithm can simultaneously process multiple images from the dataset concurrently, which does not impact the optimal solutions since the aggregate of their loss values does not affect the back-propagation process. Subsequent operations, such as sorting and hypervolume calculation for each image, can be efficiently executed in parallel using multithreading on the CPU.
IV-B Adversarial Hypervolume Training
Base Training Framework. Our adopted training framework integrates the TRADES framework [44] to serve as the cornerstone of our methodology. Employing this framework, along with a perturbation level, enables the generation of adversarial examples that augment the training process.
Ascending Perturbation Levels. By progressively elevating the perturbation levels, we replicate adversarial examples stemming from the identified adversarial frontier, in accordance with the aforementioned algorithm. This method has demonstrated efficacy in enhancing model robustness and offers greater stability than conventional adversarial training approaches [59].
Algorithm 2 Adversarial Hypervolume Training (AH Training)
1:  Input: network $f$ parameterized by $\theta$, dataset $\mathcal{D}$, total number of epochs $T$, attack epsilon step $\eta_{1}$, attack steps $K$
2:  Output: Robust network $f_{\theta}$
3:  Randomly initialize network $f_{\theta}$
4:  for $t=1$ to $T$ do
5:     $\epsilon_{t}\leftarrow t\cdot\epsilon/T$
6:     for $i=1$ to $|\mathcal{D}|$ do
7:        for $k=1$ to $K$ do
8:           $x^{\prime}_{i}\leftarrow\prod_{\mathbb{B}(x_{i},\epsilon_{t})}(x_{i}+\eta_{1}\nabla\mathcal{L}_{\text{conf}}(f,x_{i},y_{i},\epsilon))$
9:        end for
10:     end for
11:     $\mathcal{D}^{\prime}=\left\{x^{\prime}_{1},\ldots,x^{\prime}_{m}\right\}$
12:     $f_{\theta}\leftarrow\text{Train\_TRADES}(f_{\theta},\mathcal{D}\cup\mathcal{D}^{\prime})$
13:  end for
IV-C Convergence and Complexity
Assumption 2.
V Evaluation
TABLE I: Overall Results. Evaluation results of standard and adversarially trained models on the CIFAR-10 dataset [53]. Best results within the same model structure are in bold, and best overall results are highlighted with gray background.
ID Model Structure Defense Method APGD-MAR-$\ell_{2}$ Attack $\ell_{\infty}$ Attack
Accuracy (%) AH ($\text{mean}_{\text{std}}$) Accuracy (%) AH ($\text{mean}_{\text{std}}$)
Clean Robust Clean Robust
0 WideResNet-28-10 None $\mathbf{94.78}$ $\mathbf{00.35}$ $\mathbf{0.1688_{0.1636}}$ $\mathbf{94.78}$ $\mathbf{00.00}$ $\mathbf{0.0611_{0.0864}}$
1 ResNet-18 Ours $\mathbf{93.50}$ $76.28$ $\mathbf{0.7250_{0.3636}}$ $\mathbf{90.15}$ $54.70$ $\mathbf{0.4622_{0.3805}}$
2 ResNet-18 Rade et al. (2022) [60] $90.57$ $\mathbf{76.34}$ $0.5889_{0.3441}$ $86.86$ $\mathbf{57.71}$ $0.3641_{0.3201}$
3 ResNet-18 Sehwag et al. (2022) [55] $89.76$ $74.90$ $0.6382_{0.3860}$ $84.59$ $57.13$ $0.4292_{0.3745}$
4 ResNet-50 Augustin et al. (2020) [56] $\mathbf{91.08}$ $\mathbf{74.08}$ $\mathbf{0.5482_{0.3706}}$
5 WideResNet-28-10 Wang et al. (2023) [45] $\mathbf{95.16}$ $\mathbf{83.84}$ $\mathbf{0.6141_{0.2954}}$ $\mathbf{92.44}$ $\mathbf{72.35}$ $\mathbf{0.4286_{0.2970}}$
5 WideResNet-28-10 Rebuffi et al. (2021) [61] $91.79$ $79.01$ $0.5773_{0.3357}$ $87.33$ $62.37$ $0.3936_{0.3300}$
6 WideResNet-34-10 Sehwag et al. (2022) [55] $90.93$ $\mathbf{77.69}$ $\mathbf{0.6708_{0.3812}}$ $86.68$ $61.77$ $\mathbf{0.4993_{0.3972}}$
7 WideResNet-34-10 Augustin et al. (2020) [56] $\mathbf{92.23}$ $76.75$ $0.5691_{0.3654}$
8 WideResNet-34-10 Huang et al. (2021) [62] - - $90.56$ $63.55$ $0.4992_{0.3608}$
9 WideResNet-34-10 Huang et al. (2021) [62] (EMA) - - $\mathbf{91.23}$ $\mathbf{64.57}$ $0.4978_{0.3605}$
10 WideResNet-70-16 Rebuffi et al. (2021) [61] $\mathbf{92.41}$ $\mathbf{80.56}$ $\mathbf{0.6165_{0.3376}}$ $\mathbf{88.54}$ $\mathbf{66.04}$ $\mathbf{0.4395_{0.3415}}$
TABLE II: Wilcoxon rank-sum test. The best results are marked in bold and gray background. The mean value is reported, the mark (+//=+/-/=+ / - / =, better, worse, and equal) and the $p$-value are shown in the parentheses.
Defense Method APGD-MAR-$\ell_{2}$ Attack $\ell_{\infty}$ Attack
Robust Acc. AH Robust Acc. AH
Ours (Mean Value) $\mathbf{76.35}$ $\mathbf{0.7221}$ $55.36$ $\mathbf{0.4626}$
Sehwag et al. (2022) [55] $74.43(-,0.0001)$ $0.6371(-,0.0001)$ $57.32(+,0.0002)$ $0.4292(-,0.0001)$
Rade et al. (2022) [60] $76.34(=,0.5932)$ $0.5891(-,0.0001)$ $\mathbf{57.47(+,0.0002)}$ $0.3639(-,0.0001)$
V-A Experimental Setup
Datasets and Models. To evaluate adversarial robustness comprehensively, we calculate the adversarial hypervolume of models equipped with various defensive strategies and adversarial training techniques on both popular benchmark dataset CIFAR-10 dataset [53] and natural image dataset, ImageNet dataset. [63].
Datasets. For CIFAR-10 dataset [53], it consists of $50,000$ training images and $10,000$ testing images, each with dimensions of $32\times 32$ pixels in full color, across ten distinct categories. This dataset is recognized as a pivotal benchmark for adversarial robustness evaluation. Additionally, our training includes semi-supervised dataset generated by generative models from Wang et al.’s work [45]. For ImageNet dataset [63], its test set contains $100,000$ images from $1,000$ objective classes. We utilize it for providing references of our proposed metric.
Models. We utilize models from RobustBench [16], which encompasses both standardly and adversarially trained models from SOTA works in the field of adversarial robustness [45, 61, 55, 60, 56, 62, 64, 65].
Attack for Evaluation. To conduct the evaluation, a modified version of the APGD attack, named APGD-MAR, was utilized within the AutoAttack framework [14]. The APGD-MAR, an adaptation of the APGD-CE attack, is specifically designed to identify adversarial examples by employing a marginal loss function, which is a fundamental component of our metric’s implementation. Empirical evidence has demonstrated that employing solely projected gradient-based attacks is effective for assessing adversarial robustness, as observed by the negligible difference in reported adversarial accuracy when compared to the Square Attack [14].
Attack parameters were standardized across evaluations, with the number of iterations fixed at $20$ and the maximum perturbation magnitude at $8/255$ for the $\ell_{\infty}$ attack and $0.5$ for the $\ell_{2}$ attack on the CIFAR-10 dataset. For the ImageNet dataset, these parameters were adjusted to $20$ steps and a maximum perturbation of $4/255$ for the $\ell_{\infty}$ attack, , in line with the attack constraints in RobustBench [16]. Furthermore, the perturbation magnitude was divided into $N=10$ equal intervals to facilitate the computation of the adversarial hypervolume.
Adversarial Training. Adversarial training enhances the model robustness by utilizing augmented data samples crafted using various attack strategies. In this work, we use pre-trained adversarially trained models using SOTA training methods, including [1, 45, 61, 55, 60, 56, 62, 66]. These models rank highly on the RobustBench leaderboard [16] and are widely used as strong robust baselines.
V-B Overall Results
The clean accuracy, adversarial accuracy, and adversarial hypervolume of the models trained and evaluated under both $\ell_{2}$ and $\ell_{\infty}$ threat models on the CIFAR-10 dataset are presented in TABLE I. Specially, the mean and the standard deviation of the adversarial hypervolume metrics calculated on the test set are reported. Generally, the adversarial accuracy and adversarial hypervolume of the models are consistent, with the adversarial hypervolume providing a more comprehensive evaluation of the model’s robustness. In the following sections, we analyze the results and discuss the implications of the findings.
Effectiveness of AH Training. The AH training algorithm has demonstrated substantial efficacy in augmenting the adversarial robustness of machine learning models, as evidenced by the improved adversarial accuracy and hypervolume metrics. Specifically, the models utilizing our training algorithm achieved the best adversarial hypervolume among the models trained with the same structure under both $\ell_{2}$ and $\ell_{\infty}$ attacks. Moreover, our models reported comparable adversarial accuracy to the SOTA models out of all model structures, demonstrating the effectiveness of our training algorithm.
Standardly v.s. Adversarially Trained. The adversarial hypervolume of the standardly trained models is significantly lower than that of the adversarially trained models as shown in TABLE I. As the initial confidence score of the standardly trained models is usually high, we can conclude that their confidence score decreases significantly when exposed to adversarial examples with minimal perturbation, leading to a lower adversarial hypervolume. It is evidenced by the example in Figure 2.
AH for Model Selection. When two models have similar adversarial accuracy, adversarial hypervolume serves as an additional metric to differentiate their robustness, as shown in the comparison between our model and the model trained by Rade et al. [60]. Adversarial hypervolume provides a more comprehensive evaluation of the model’s robustness, which is crucial for selecting the most robust model for deployment.
V-C Further Validation of Effectiveness
To further validate the effectiveness of our method and diminish the potential impact of randomness, we conducted a Wilcoxon rank-sum test on our model and the model trained by Rade et al. [60] and Sehwag et al. (2022) [55] under $\ell_{2}$ and $\ell_{\infty}$ attacks. Specifically, we adopted the same training and evaluation settings for our model in Section V-B and computed the adversarial accuracies and adversarial hypervolumes of the models under $10$ different random seeds. The two baseline models are from from RobustBench [16]. For fair comparison, both the baseline models and our model are ResNet-18 models.
Results The results of the Wilcoxon rank-sum test are shown in Table II. The table contains the results robust accuracy and adversarial hypervolume under $\ell_{2}$ and $\ell_{\infty}$ attacks, with the comparison results determined with the significance level of $p=0.05$. Our methods achieved the highest adversarial hypervolume under both attacks, with the highest robust accuracy under the $\ell_{2}$ attack. For the robust accuracy under the $\ell_{\infty}$ attack, our model achieved comparable performance with only $2\%$ lower than the best model. The Wilcoxon rank-sum test results further confirm the effectiveness of our method in enhancing the adversarial robustness of the models.
TABLE III: Ablation Study. Ablation study for our training algorithm under CIFAR-10 dataset.
Defense Method APGD-MAR-$\ell_{2}$ Attack $\ell_{\infty}$ Attack
Accuracy (%) AH ($\text{mean}_{\text{std}}$) Accuracy (%) AH ($\text{mean}_{\text{std}}$)
Clean Robust Clean Robust
None $\mathbf{93.50}$ $\mathbf{76.28}$ $0.7250_{0.3636}$ $\mathbf{90.15}$ $\mathbf{54.70}$ $0.4622_{0.3805}$
w./o. ascending $93.59$ $74.12$ $0.6978_{0.3713}$ $89.84$ $53.03$ $0.4556_{0.3809}$
w./o. data $90.49$ $63.74$ $\mathbf{0.7292_{0.3765}}$ $86.60$ $40.84$ $\mathbf{0.5626_{0.4019}}$
TABLE IV: Input Transformation. Results on Input Transformation Defense Methods. Clean is for clean accuracy and Robust is for robust accuracy. The best results are marked in bold and gray background.
Method Clean APGD-MAR-$\ell_{2}$ Attack $\ell_{\infty}$ Attack
Robust AH Robust AH
None $\mathbf{94.78}$ $00.35$ $0.1688_{0.1636}$ $00.00$ $0.0611_{0.0864}$
JPEG [39] $91.28$ $07.61$ $0.0521_{0.2047}$ $03.19$ $0.0001_{0.0098}$
FS [19] $94.24$ $03.28$ $0.0418_{0.1989}$ $01.61$ $0.0394_{0.1945}$
SPS [19] $87.33$ $\mathbf{20.21}$ $\mathbf{0.2033_{0.3765}}$ $\mathbf{05.91}$ $\mathbf{0.0105_{0.0929}}$
V-D Results on Input Transformation
Results. Table IV presents the results for the various input transformation methods. All methods display significantly lower adversarial accuracies under both attack types. Furthermore, when employing prevalent perturbation budgets, the $\ell_{\infty}$ attack is observed to be more potent than the $\ell_{2}$ attack, as evidenced by the lower adversarial accuracy and hypervolume. Of all the input transformation methods evaluated, only spatial smoothing imparts some resistance to $\ell_{2}$ attack, whereas none effectively withstand $\ell_{\infty}$ attack. The adversarial hypervolume metric suggests that while these transformations can resist minor noise, they are ineffective against substantial perturbations.
V-E Ablation Study
The efficacy of our adversarial hypervolume training techniques was further investigated through an ablation study. The study involved two modifications:
Results. The findings, as presented in TABLE III, indicate that the standard approach outperforms alternative strategies concerning adversarial robustness, achieving the highest performance in adversarial accuracy during both $\ell_{2}$ and $\ell_{\infty}$ attacks, with the slight expense of clean accuracy. Elimination of the ascending technique resulted in a marginal reduction in robustness metrics.
Mismatch. In TABLE III, we noticed an unusual increase in the adversarial hypervolume when the synthetic data was removed, along with a decrease in the adversarial accuracy. This discrepancy suggests that the model can be over-confident on its predictions when the synthetic data is removed, leading to a lower adversarial accuracy but a higher adversarial hypervolume. This phenomenon reveals that both adversarial accuracy and adversarial hypervolume are essential when considering the evaluation of adversarial robustness.
Further Results on Ascending. The effectiveness of ascending strategy of the perturbation budget is further validated, with the results under $\ell_{\infty}$ attack shown in Figure 4 and TABLE V. The ascending strategy significantly improves the adversarial hypervolume of the models under $\ell_{\infty}$ attacks, as evidenced by the higher adversarial hypervolume of the models trained with the ascending strategy compared to the models trained with the fixed budget strategy.
(a) Clean Accuracy
(b) Robust Accuracy
Figure 4: Accuracy Comparison. Results of the clean and robust accuracy trained under fixed budget and ascending budget strategy.
TABLE V: Ascending Strategy. Adversarial hypervolume metric of models trained using different perturbation budgets.
$\epsilon$ None No ascending
$4/255$ $0.5444_{0.3686}$ $\mathbf{0.5893_{0.3699}}$
$8/255$ $\mathbf{0.4656_{0.3805}}$ $0.4622_{0.3809}$
$12/255$ $\mathbf{0.3431_{0.3429}}$ $0.3288_{0.3421}$
$16/255$ $\mathbf{0.2319_{0.2820}}$ $0.1978_{0.2585}$
V-F Results of ImageNet Dataset
As a further validation of our proposed metric, we conducted experiments on the ImageNet dataset to evaluate the adversarial robustness of models under the $\ell_{\infty}$ attack. The results are shown in Table VI. Generally, the adversarial accuracy and adversarial hypervolume of the models are consistent, both indicating a increase in adversarial robustness with advanced defense methods and more complicated model structures. Further investigations on the ImageNet dataset will be conducted in our future work to provide a more comprehensive understanding of the relationship between adversarial accuracy and adversarial hypervolume.
TABLE VI: Results on ImageNet dataset. Results on ImageNet dataset under $\ell_{\infty}$ Attack.
Model Structure Defense Method Accuracy AH
Clean Robust
ResNet-50 Standard $\mathbf{75.89}$ $26.98$ $0.3093_{0.3592}$
ResNet-50 Wong et al. [64] $52.71$ $49.55$ $0.3100_{0.3905}$
ResNet-50 Engstrom et al. [66] $61.93$ $58.84$ $0.3718_{0.4038}$
ResNet-50 Salman et al. [65] $63.32$ $\mathbf{60.61}$ $\mathbf{0.3721_{0.3985}}$
ResNet-18 Salman et al. [65] $51.92$ $49.00$ $0.2424_{0.3449}$
V-G Convergence Results
VI Discussion
VI-A Why Adversarial Hypervolume?
VI-B Intuition of the Proposed Metric
VI-C Extension to Other Types of Attacks
The adversarial hypervolume metric is computed under a bi-objective adversarial problem setting, where the perturbation budget and confidence score are the two objectives. Therefore, the extension of the metric can be divided into two directions: new implementation and new problem definition.
New Implementation. The adversarial hypervolume metric can be directly implemented in $\ell_{0}$ attacks, with the perturbation budget being the $\ell_{0}$ norm and the confidence score being the marginal confidence score. Recently, $\ell_{0}$ adversarial attacks, also known as sparse adversarial attacks, have attracted increasing attention [47], thereby necessitating the development of a robust metric for evaluating the adversarial robustness of models against $\ell_{0}$ attacks. Moreover, under this attack setting, the problem definition can be extended to a tri-objective problem by incorporating the perturbation budget as an additional objective, leading to a more comprehensive evaluation of the model’s robustness. We plan to explore this direction in our future work.
New Problem Definition. As new types of adversarial attacks emerge, e.g. adversarial patch attacks [68], we can extend the adversarial hypervolume metric to accommodate these new attack settings. For instance, in adversarial patch attacks, the perturbation budget can be defined as the size of the patch, while the confidence score can be the marginal confidence score. By incorporating these two objectives, we can evaluate the robustness of models against adversarial patch attacks. We aim to explore this direction in our future work.
VI-D Connection and Difference with Other Metrics